# Lexical categories of stem-forming roots in Mapudüngun verb forms

### v2

@@ 4.2 Tables
c. In reality, to convey the meaning of ‘to relate’, the compound che-m-ye- needs to be formed, which is analysed as -NN.che_person_people+CA.m34+TV.ye_to_carry. This is the way to express a relationship with someone in Mapudüngun. For example, laku-ye-fi-ñ translates to ‘he is my paternal grandfather’, which literally means ‘I carry him as a grandfather’ (Becerra, 2011: 115).
@@
d. In the case of interrogatives like ‘‘what, which, how’’, etc., staring with chu-, these are formed by adding other suffixes. For example, chu-m-al means ‘what for’, chu-m-nge-chi means ‘how’, chu-m-ül means ‘when’, and so on.
@@
e. It is very likely that neykü is a lexicalised form, composed of the root nel- listed in line 39, and the unproductive suffix -kü-, which is found in a few verbal roots but has not been documented or described in existing literature.
f. pelo is a lexicalised form of a nominalised verb (see example E.28.a in Chandía, 2025).
@@
g. The form püra-m- generates compounds with many action verbs to indicate an upward direction, such as in tralkatunpüramün, which means ‘to shoot into the air’. It is also used figuratively, as in kimpüramün, meaning ‘to become aware’; literally: ‘to raise knowledge’.
h. When pütre is nominalised as pütrem -AJ.pütre_blazing+CA.m34+NOM.ø, it means ‘cigar’ or ‘tobacco’.
@@
i. trongum- is to thatch a ruka (a traditional Mapuche house) with reed or tall grass, these materials must be placed very densely, ensuring a good level of thickness.

### v3

@@ 4.2   Tables
c In reality, to convey the meaning of ‘to relate’, the compound che-m-ye needs to be formed, which is analysed as -nn.che_person_people+ca.m+tv.ye_to-carry. This is the way to express a relationship with someone in Mapudüngun. For example, laku-ye-fi-ñ translates to ‘he is my paternal grandfather’, which literally means ‘I carry him as a grandfather’ (Becerra, 2011: 115).
@@
d In the case of interrogatives like “what, which, how”, etc., starting with chu-, these are formed by adding other suffixes. For example, chu-m-al means ‘what for’, chu-m-nge-chi means ‘how’, chu-m-ül means ‘when’, and so on.
@@
e It is very likely that neykü is a lexicalised form, composed of the root nel- listed in line 39, and the unproductive suffix -kü, which is found in a few verbal roots but has not been documented or described in existing literature.
f pelo is a lexicalised nominalised verb, likely from pe- ‘sight, vision’ + fossilised suffix -lo (no longer productive), meaning ‘light, clarity’. For the origin of similar lexicalised forms, see footnote 18. For an example of pelo in a verbal form, see the companion valency article (Chandía, 2026, E.28a).
@@
g The form püra-m generates compounds with many action verbs to indicate an upward direction, such as in tralkatunpüramün, which means ‘to shoot into the air’. It is also used figuratively, as in kimpüramün, meaning ‘to become aware’; literally: ‘to raise knowledge’.
h When pütre is nominalised as pütrem (-aj.pütre_blazing+ca.m+nom.ø), it means ‘cigar’ or ‘tobacco’.
@@
i trongum- is to thatch a ruka (a traditional Mapuche house) with reed or tall grass; these materials must be placed very densely, ensuring a good level of thickness.
?refname?
  • A. Álvarez, R. Aranovic, R. Brown, J. Carbonell, C. Fasola, A. Lavie, L. Levin, A. Font-Llitjos, C. Monson, E. Peterson, K. Probst, and R. Vega (2005) Informe final del Proyecto Avenue / Mapudungun: Desarrollo de herramientas informáticas para el mapudüngun que se habla en Chile. Technical report Language Technologies Insitute. Carnegie Mellon University, Pensilvania, EE. UU.. Cited by: §1.1.2.
  • R. Becerra (2011) Una propuesta funcional y tipológica de descriptores sintácticos para el mapudungun. Alpha (Osorno), pp. 111–125 (es). Note: PDF External Links: ISSN 0718-2201, Link Cited by: §4.2.
  • A. Chandía (2012) Düngupeyem_alfa_v0.1: un prototipo de analizador morfológico para el mapudüngun a través de transductores de estados finitos. Máster en Ciencia Cognitiva y Lingüística, Universidad de Barcelona. Note: PDF Cited by: §1.
  • A. Chandía (2014) CORLEXIM Corpus lexicográfico del mapudüngun. Note: Online resource http://www.chandia.net/corleximAvailable dictionaries: 1. Valdivia, L. de (1606): mapudüngun-castellano. 2. Febrés, A. (1765): castellano-mapudüngun. 3. Febrés, A. ([1765] 1882): mapudüngun-castellano. 4. Febrés, A. (1846a): mapudüngun-castellano. 5. Febrés, A. (1846b): castellano-mapudüngun. 6. Augusta, F. de (1916b): castellano-mapudüngun. 7. Augusta, F. de (1916a): mapudüngun-castellano. 8. Smeets, I. (2008): mapudüngun-inglés Cited by: item 1, item 2, item 3, item 4, §1.1.3, §1.1.3, §1.1.3, §1.1.3, §1.1.3, §2.1, footnote 5.
  • A. Chandía (2021) FST Morphological Analyser and Generator for Mapudüngun. arXiv. Note: PDF External Links: Document, 2109.09176 Cited by: §1, §1, footnote 4.
  • A. Chandía (2022) A Mapudüngun FST Morphological Analyser and its Web Interface. In Proceedings of the Language Resources and Evaluation Conference, Marseille, France, pp. 6540–6547. Note: PDF Cited by: §1.
  • A. Chandía (2026) Valency classification of mapudungun verbal roots. established by the language’s own morphotactics. External Links: 2604.00789, Link Cited by: §1.2, §1.2, §1, §1, item 6., §2.2, §2.3, §2.5, §2, §3, §3, §4.2, §4.2.
  • F. de Augusta (1903) Gramática araucana. Imprenta Central, J. Lampert, Valdivia. Note: PDF Cited by: item d..
  • L. Golluscio (2007) Morphological causatives and split intransitivity in mapudungun. International Journal of American Linguistics 73 (2), pp. 209–238. External Links: ISSN 00207071, 15457001, Link Cited by: §2.5, §2.5, §4.2.
  • E. Longkon (2011) Morfología y aspectos del mapudüngun. Biblioteca de signos, Universidad Autónoma Metropolitana, México. External Links: ISBN 9786074775532 Cited by: §2.5.
  • C. Martinez and M. Carmona (2023) El Vocabulario de Bernard Havestad (1777 – Pars Quarta), versión mapudungun-español. . Note: Online resource https://benmolineaux.github.io/bookshelf/ Cited by: footnote 6.
  • I. Smeets (2008) A Grammar of Mapuche. De Gruyter Mouton, Berlin, Boston. Note: PDF External Links: Document, ISBN 9783110211795 Cited by: §1, item a., item b., item c., 1st item, §1.2, §1, §1, §1, §1, §1, §2.1, §2.5, §2.5, §3, §4.2, footnote 2.
  • B. Villena (2016) Innovación léxica en mapudungún: genuinidad, productividad y planificación. Doctoral Dissertation, Universitat Pompeu Fabra. External Links: Link Cited by: §1.1.3, §1.1.3, §1.1.3, §1.1.3, footnote 12.
  • A. Wilhelm, Pascual. Koña, and R. Lenz (1930) Vida y costumbres de los indígenas araucanos en la segunda mitad del siglo XIX. Imprenta Universitaria, Santiago, Chile. Note: PDF External Links: LCCN 37025221 Cited by: 2nd item, §1.2, §2.1, §2.3, §2.5, §3, §4.2.
  • F. Zúñiga (2006) Mapudüngun. el habla mapuche. Centro de Estudios Públicos, Santiago, Chile. Note: PDF External Links: ISBN 9783895869761 Cited by: footnote 10.